# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems for Social Media

### v2

@@ 7.1. Participation
We define participation as the design choices that shape who uses the system and what different users are allowed to do. Table 2 shows all of the participation dimensions and options. Choices that platforms make in this aspect range from those that encourage user participation in the system, to those that discourage it.
@@
The most basic dimension is, “Who sees published Notes?” The options that platforms choose as answers to this question will greatly affect the system’s impact. The most participatory implementations like X’s allows all platform users to see these notes, while less participatory implementations limit the audience in some way beyond a user’s control, for example by their location. An even less participatory option would be to make notes an optional feature that users must opt in to see, which would greatly limit the system’s impact.
@@
Other dimensions of participation determine who can contribute to the system, such as: “Who are the context contributors?” These dimensions shape the quality, quantity, and variety of the notes that the system produces. All implementations we examined rely on platform users as contributors, but we see some variation in what these users are allowed to contribute. For example, YouTube currently does not allow users to provide ratings; instead, those are provided by third-parties who are paid by the platform. X allows AI systems to contribute notes via an API. All of the implementations impose eligibility restrictions on users, ensuring that platforms maintain discretion over who is admitted as a contributor. Yet we can imagine a system that would automatically admit all eligible applicants, or even one with no eligibility restrictions, where all users would be contributors by default.
@@
Besides decisions that limit the set of possible contributors, it is also important to ask questions about how platforms encourage these contributors to participate, such as: “How are contributors incentivized?” While it would be possible for platforms to pay contributors, all of the implementations that we examined rely on volunteer labor to some extent. These approaches may choose to incentivize contributions through reputation systems, for example by publicly associating contributors with either their user accounts or with pseudonyms. There is also the important question of: “How are notes and ratings solicited?” All platforms allowed contributors to add notes to posts that they organically come across while using the platform. However, this design choice could result in skewed post selection if, for example, contributors follow accounts from one end of the ideological spectrum more than the other. Some platforms may also take more active approaches that solicit notes for specific kinds of posts, or ratings for specific suggested notes, either through in-app feeds or notifications, as shown in Table 2. These approaches could be used to counter bias in the system, or to encourage notes on posts that are most likely to do harm, such as those that are going viral.
@@ 7.2. Inputs
Free-text inputs
@@
A fixed set of options (i.e., helpful, somewhat helpful, not helpful) to select from.
@@
Table 3. Inputs dimensions and options, as well as platforms that have made this design choice.:
@@
Contributions are shaped by the affordances that platforms provide to contributors to create their input, including the user interfaces for producing notes and ratings. These interfaces may allow free-form, or structured, input. Such decisions may have an effect on what information contributions contain and what they look like. For example, Meta’s implementations have a required “source URL” field, which based on the findings mentioned in Section 6, may lead to more helpful notes. The content of contributions is also shaped by additional restrictions placed on them by platforms. For example, all of the implementations that we examined required contributions to adhere to the platform’s content policy. Finally, design choices that require contributors select from a set of fixed options, may be both easier to use and produce more consistent inputs. In these ways, design decisions in this aspect will affect the volume and quality of both notes and ratings.
A final interesting question is, “Can the published note on a post change over time?”. Some implementations will give contributors the ability to edit or delete their notes, while others will prevent this. Additionally, the published note on a post can change if the curation process that selects notes (discussed below) is re-run over time. To deal with this, X’s implementation added a feature that “locks” the published note after a certain amount of time, preventing it from changing.
@@ 7.3. Curation
Identify notes that appeals across the ideological spectrum (“bridging”).
@@
Table 4. Curation dimensions and options, as well as platforms that have made this design choice.:
We define curation as the design choices that affect how notes are selected to be displayed. Table 4 shows all of the associated dimensions and options. Design decisions here affect the quality of the notes that are displayed, the potential for bias in the note selection, as well as the speed and volume at which they appear.
Perhaps the most important design question is “What process determines which notes are shown alongside posts?” The existing systems that have recently grown in popularity all involve algorithmic decisions by bridging-based algorithms that, broadly speaking, identify notes that have been rated as “helpful” by contributors who have disagreed in their past ratings. Even though these bridging algorithms are currently popular, alternatives to this scheme are possible: for example, Twitter’s first implementation of Birdwatch used a majority voting algorithm. Alternatively, it could be possible to achieve bridging goals using contributor information other than their past ratings. Algorithmic approaches are also not the only option, as one could imagine a CCS that selects notes via editorial decisions by either the platform or contributors. Another possible approach is to select notes via digital juries (fanDigitalJuriesCivicsOriented2020) of users, which could be selected either randomly or based on other criteria.
A related question is, “What (alleged) control does the platform have over what is displayed?” Platforms can exert varying levels of control over the output of these systems, although current implementations all intentionally minimize the platform’s role. Still, all platforms do retain control to make editorial decisions in the case of content policy violations.
@@ 7.4. Presentation
How are a note’s ratings shown on a post?
Only whether or not it is “helpful”.
Some indicator of how helpful it is.
Table 5. Presentation dimensions and options, as well as platforms that have made this design choice.:
We define presentation as the design choices about how published notes are displayed to users. Table 5 shows all of the associated dimensions and options. Design decisions here affect the impact of a published note on users who see the associated post.
The biggest question is, “Where are published notes presented to users?” Answers to this question affect whether or not users will ever see published notes. Note presentation can range from aggressive to subtle: notes can be mandatory overlays that users must dismiss to see the associated post; notes can be shown alongside posts where they can be easily ignored; or platforms may simply indicate that a post has a note and require users to click to view them. Additionally, there is the question of which instances of the post have the published note displayed: just the original, or also re-shares of the post that appear across the platform. This will affect the number of users who have the potential of being informed by a published note. Platforms must also decide: “Who is notified about notes when they are published?” Notifications here can be used to retroactively influence the understanding of users who saw a post before it had a published note.
Other important questions affect the amount of information about notes that is conveyed to users. Twitter’s original implementation of Birdwatch showed several “helpful” notes at a time, but all of the current implementations that we examined only show one. Relatedly, platforms may provide details about a note’s author or ratings, which could inform how a user interprets a note.
@@ 7.5. Platform Treatment
How are posts with notes on them treated by the platform?
@@
Table 6. Platform Treatment dimensions and options, as well as platforms that have made this design choice.:
We define Platform Treatment as the design choices about how the platform treats posts that have published notes. Table 6 shows all of the associated dimensions and options. Design decisions here affect the impact of CCS systems on the spread of posts.
The central question is, “How are posts with notes on them treated by the platform?” All current implementations make a point of emphasizing that they do not algorithmically penalize posts with notes on them (for example, by down-ranking in recommendation algorithms). But this need not be the case: perhaps certain types of notes provide a useful signal about post quality and platforms may want to suppress them. Alternatively, the platform may impose restrictions on posts with notes, like prohibiting them from being monetized or from displaying next to ads.
It is also important to ask “How do notes interact with other content moderation features?” Some platforms have shut down their fact-checking programs and replaced them with CCS. But, all of the platforms that we surveyed still have content moderation systems of some kind, which a CCS complements. Finally, YouTube and TikTok display multiple “information reliability” indicators alongside posts in tandem with notes, such as platform-provided banners or labels.
@@ 7.6. Transparency
Table 7. Transparency dimensions and options, as well as platforms that have made this design choice.:
We define transparency as the design choices that influence how the public can observe and evaluate the functioning of the system. Table 7 shows all of the associated dimensions and options. Design decisions here may affect the perception of the CCS by users, the public, and the research community.
Of central interest is, “How public is the note curation system?” The system’s source code and data can be public in order to allow outside study. In addition to being public, platforms could take additional steps to make their source code and data verifiable, or to allow outsiders to verify that the public code and data is indeed the same code and data that is powering the platform’s implementation. Or, platforms could do none of this and instead keep everything private, which is the approach chosen by all platforms except for X.
To fully evaluate the efficacy of such a system, we also need to ask, “How public is the data needed to evaluate the impact of notes on engagement?” This data would allow the public to holistically evaluate the CCS as opposed to other systems of content moderation. Currently this data is very hard to access, which makes holistically assessing the efficacy of these systems a challenge.
@@ 8. Discussion
Through our systematic literature review and analysis of real-world CCS implementations, we have produced a theoretical framework that can help the HCI research community better reason about, and design, these systems. In this section, we discuss future directions for CCS research in HCI. Specifically, we focus on how CCS fit within the broader field of content moderation, highlight underexplored portions of the design space, and reflect on the normative implications of CCS design choices.
@@ 8.1. Can CCS Replace Content Moderation?
CCS were created with the stated goal of addressing the misinformation problem (IntroducingBirdwatchCommunitybaseda), but it is unclear if the platforms implementing them are serious about this aim. For example, many see Meta’s choice to replace their U.S. fact-checking program with a CCS (heatheraMoreSpeechFewer2025) as a thinly-veiled effort to curry political favor with a new presidential administration that is critical of fact-checking (nixMetaEndsFactchecking2025). In fact, from a platform’s perspective, CCS have other appealing properties: they cost less than professional fact-checkers, they shield platforms from accusations of biased policy enforcement, and they offload the difficult responsibility of distinguishing fact from fiction onto users. It may be expedient for platforms to build CCS in the current moment, but it is unclear if such systems will survive if the political climate shifts. Still, as long as CCS are in a position to exert such influence over the information ecosystem, it is important for the HCI research community to study them. Most importantly, we emphasize the need to better understand how CCS perform compared to other content moderation systems. CCS may perform well in certain cases, but if on the whole they perform worse than the systems they are replacing, for example on the question of coverage, they should be used thoughtfully.
While it has been suggested that they can, CCS cannot effectively replace a fuller spectrum of content moderation tools, and were not designed to do so (elliottElonMusksMain). Content moderation is a broad term that encompasses a variety of different goals, including preventing crimes, mitigating lawful but awful behavior, and promoting high quality content (grimmelmann2015virtues). The tools used as a part of content moderation are similarly varied and include both “hard” approaches, like content take-downs and account bans, and “soft” approaches, such as content labeling (zannettouWonElectionEmpirical2021). Systems that provide context are narrow in their their goal and mechanism: they are a soft moderation attempt to aid content interpretation by providing users with additional information (matamoros25). And yet, the empirical results summarized in this paper demonstrate that CCS may play a meaningful role in the content moderation ecosystem, and it is worth exploring how they may integrate with other parts of it. For example: how might CCS be integrated into platforms’ automated moderation systems? In the Platform Treatment design aspect, none of the implementations that we examined used the CCS output to affect the visibility of posts. Perhaps these systems can be integrated into platform content recommendations algorithms to promote or demote posts based on their notes.
@@ 8.2. Underexplored CCS Design Aspects
Our design space reveals several ways forward for a HCI research agenda on CCS. Several aspects are underexplored by current research, particularly the Participation aspect. Encouraging positive participation in collaborative computing systems is a classic topic in HCI research (krautBuildingSuccessfulOnline2012) and one in which the community could make strong contributions. One dimension of participation which has not yet been explored is how platforms may play a more active role in soliciting notes and ratings from contributors. Past HCI work on how to best make use of the various skills of crowds may be instructive here (bernsteinSoylentWordProcessor2015; kitturFutureCrowdWork2013). It is worth noting that the more consequential these systems become, the more they will be targets for negative forms of participation as well, primarily gaming (danielQualityControlCrowdsourcing2018; epsteinWillCrowdGame2020). Bridging algorithms are designed to be resistant to certain types of gaming (wojcikBirdwatchCrowdWisdom2022b), but vulnerabilities may emerge. One other area for study is the Presentation aspect, specifically how different approaches to visually displaying notes may affect user understanding of misleading claims (gamageLabelingSyntheticContent2025; martelMisinformationWarningLabels2023). Another promising area is the Inputs aspect, where recent work has begun to explore the ways in which LLMs can supplement human effort and increase the quality and quantity of notes (supernotes; mohammadiAIFeedbackEnhances2025a). Of particular interest is X’s recent decision to allow AI note contributors via a new API (liScalingHumanJudgment2025a). How such a decision affects the quantity and quality of notes on the platform should be a near-term focus of HCI research. Of course, studying how changes to CCS impact information dynamics across a platform will be difficult to study without more innovations in the Transparency aspect. X’s approach to transparency in their Community Notes system has been a boon for studying CCS. Hopefully other platforms will soon adopt similar approaches to support ongoing studies of CCS effectiveness.
@@ 8.3. Normative Guidelines for CCS Design
CCS are complex sociotechnical systems (chernsPrinciplesSociotechnicalDesign1976) that integrate people, algorithms, and speech. These systems’ growing influence over the information shown on major social media platforms means that the design choices made in their implementation can have significant social consequences (friedmanBiasComputerSystems1996; flanaganEmbodyingValuesTechnology2008). With such high stakes, it is crucial that a human-centered research agenda consider the implications of different design choices, in order to holistically evaluate these systems’ impacts. One way to do so is to identify the primary normative concerns of such a system. We emphasize three interconnected concerns that should guide CCS design: user informedness, distribution of power, and fairness.

### v3

@@ 7.1. Participation
We define Participation as the design choices that shape who uses the system and what different users are allowed to do. Table 2 shows the participation dimensions and options derived from our analysis. Platform choices in this aspect range from those that encourage user participation in the system to those that discourage it.
@@
The most basic dimension is, “Who sees published Notes?” Platforms’ answers to this question greatly affect the system’s impact. The most participatory implementations, like X’s, allow all users to see these notes, while less participatory implementations limit the audience, for example by location. An even less participatory option is making notes an optional feature that users must opt into, which would greatly limit system impact.
@@
Other dimensions of participation determine who can contribute to the system, such as: “Who are the context contributors?” These dimensions shape the quality, quantity, and variety of the notes the system produces. All implementations we examined rely on platform users as contributors, but we see variation in what these users are allowed to contribute. For example, YouTube does not allow users to provide ratings; instead, ratings are provided by third parties paid by the platform. X allows AI systems, in addition to users, to contribute notes via an API. All implementations impose eligibility restrictions on users, limiting who can be a contributor. Yet platforms also decide which eligible applicants are admitted. A more participatory system would automatically admit all eligible applicants, or could eliminate eligibility restrictions and make all users contributors by default.
@@
Beyond decisions that limit the set of possible contributors, platforms must also consider how to encourage contributor participation: “How are contributors incentivized?” While platforms could pay contributors, all implementations we examined rely on volunteer labor to some extent. These approaches may incentivize contributions through reputation systems, for example by publicly associating contributors with their user accounts or with pseudonyms. Another important question is: “How are notes and ratings solicited?” All platforms allow contributors to add notes to posts they organically encounter. However, this design choice may bias which posts receive notes if, for example, contributors only follow accounts from one end of the ideological spectrum. Section 7.3 explores ways to reduce bias when selecting which notes to publish, but the choices here affect the supply of notes available for selection. Some platforms take more active approaches and use in-app feeds or notifications to solicit notes for specific posts or ratings for specific suggested notes, as shown in Table 2. These approaches can counter supply-side bias or encourage notes on posts most likely to cause harm, such as those going viral.
@@ 7.2. Inputs
Free-text inputs.
@@
A fixed set of rating options (i.e., helpful, somewhat helpful, not helpful).
@@
Table 3. Inputs dimensions and options, as well as platforms that documented these design choices.
Inputs dimensions covering what tools are provided to contributors, which posts can have notes, restrictions on notes and ratings, and whether notes can change over time. Shows the various input methods and constraints across different platforms.
@@
Contributions are shaped by the affordances that platforms provide contributors, including the user interfaces for producing notes and ratings. These interfaces may allow free-form or structured input. Such differences may affect what information contributions contain and how they appear. For example, Meta’s implementation has a required “source URL” field, which, based on findings from Section 6, may lead to more helpful notes. Contributions are also shaped by additional platform restrictions. For example, the implementations we examined required that contributions adhere to the platforms’ content policies. Finally, design choices that give contributors fixed options may both be easier to use and produce more consistent contributions. In these ways, design decisions in this aspect affect the volume and quality of both notes and ratings.
A final interesting question is, “Can the published note on a post change over time?” Some implementations allow contributors to edit or delete their notes, while others do not. Additionally, the published note on a post can change if the curation process that selects notes (discussed below) is re-run over time. To deal with this, X’s implementation has a feature that “locks” the published note after a certain amount of time, preventing it from changing.
@@ 7.3. Curation
Identify notes that appeal across the ideological spectrum (“bridging”).
@@
Table 4. Curation dimensions and options, as well as platforms that documented these design choices.
Curation dimensions focusing on how notes are selected for display, including algorithmic decision processes, platform control levels, data sources used, and aims of algorithmic decisions. Primarily shows bridging-based approaches across platforms.
We define Curation as the design choices that affect how notes are selected for display. Table 4 shows all of the associated dimensions and options. Design decisions here affect the quality of the displayed notes, the potential for bias in note selection, as well as the speed and volume at which notes appear.
Perhaps the most important design question is “What process determines which notes are shown alongside posts?” The implementations that we studied all use bridging algorithms to, broadly speaking, identify notes that have been rated as “helpful” by contributors who have disagreed in their past ratings. While these bridging algorithms are currently popular, alternatives are possible: for example, Twitter’s first implementation of Birdwatch used a majority voting algorithm. Alternatively, bridging goals could be achieved using contributor information other than their past ratings, such as the accounts that they follow. Algorithmic approaches are also not the only option, as CCS could select notes via editorial decisions by either the platform or contributors. Another possible approach is selecting notes via digital juries (Fan and Zhang, 2020) of users, who could be selected either randomly or based on other criteria.
A related question is: “What (alleged) control does the platform have over what is displayed?” Platforms can exert varying levels of control over system output, although current implementations all intentionally minimize the platform’s role. Still, platforms retain control over editorial decisions in cases of content policy violations.
@@ 7.4. Presentation
What information about a note’s ratings are shown on a post?
Only whether the note is rated “helpful”.
Some indicator of how helpful the note is.
Table 5. Presentation dimensions and options, as well as platforms that documented these design choices.
Presentation dimensions covering where notes are displayed, who gets notified, how many notes are shown, and how authorship and ratings are presented. Most platforms show notes below posts with anonymous authorship.
We define Presentation as the design choices affecting how published notes are displayed to users. Table 5 shows all associated dimensions and options. Design decisions here affect the impact of published notes on users who see the associated posts.
The biggest question is: “Where are published notes presented to users?” Answers to this question affect whether users will ever see published notes. Note presentation can range from aggressive to subtle: notes can be mandatory overlays that users must dismiss to see the associated post; notes can be shown alongside posts where they can be easily ignored; or platforms may simply indicate that a note exists and require users to click to view it. Additionally, there is the question of which instances of the post display the published note: just the original, or also re-shares that appear across the platform. This affects the number of users who see a published note. Platforms must also decide: “Who is notified about notes when they are published?” Notifications can retroactively influence the understanding of users who saw a post before it had a published note.
Other important questions affect the amount of information about notes conveyed to users. Twitter’s original Birdwatch implementation showed several “helpful” notes at a time, but all implementations we examined show only one. Relatedly, platforms may provide details about a note’s author or ratings, which could inform how users interpret notes.
@@ 7.5. Platform Treatment
How are posts with notes treated by the platform?
@@
Table 6. Platform Treatment dimensions and options, as well as platforms that documented these design choices.
Platform Treatment dimensions showing how posts with notes are treated by platforms, including monetization restrictions, interaction with other moderation features, and relationship to fact-checking programs.
We define Platform Treatment as the design choices about how the platform treats posts that have published notes. Table 6 shows all associated dimensions and options. Design decisions here affect the impact of CCS on post spread.
The central question is: “How are posts with notes treated by the platform?” All current implementations emphasize that they do not algorithmically penalize posts with notes (for example, by down-ranking them in recommendation algorithms). But this need not be the case: perhaps certain note types provide a signal about low post quality and platforms may want to suppress associated posts. Alternatively, platform may impose restrictions on posts with notes, such as prohibiting monetization or display next to ads.
It is also important to ask: “How do notes interact with other content moderation features?” Some platforms have shut down their fact-checking programs and replaced them with CCS. However, all platforms we surveyed still have content moderation systems of some kind, which CCS complement. Finally, YouTube and TikTok display multiple “information reliability” indicators alongside posts in tandem with notes, such as platform-provided banners or labels.
@@ 7.6. Transparency
Table 7. Transparency dimensions and options, as well as platforms that documented these design choices.
Transparency dimensions covering public access to curation systems, contributor admission processes, engagement impact data, and documentation availability. Shows X/Twitter as most transparent with open source and open data approaches, while other platforms are mostly private.
We define Transparency as the design choices that influence how the public can observe and evaluate system functioning. Table 7 shows all associated dimensions and options. Design decisions here may affect the perception of CCS by users, the public, and the research community.
Of central interest is: “How public is the note curation system?” CCS source code and data can be public to allow outside study. In addition to being public, source code and data can be made verifiable, allowing outsiders to verify that the code and data indeed powers the platform’s implementation. Alternatively, platforms could keep everything private, as all platforms except X currently do.
To fully evaluate the efficacy of such a system, we also need to ask: “How public is the data needed to evaluate the impact of notes on engagement?” This data would allow the public to holistically evaluate CCS against other content moderation systems. Currently, this data is very difficult to access, making holistic assessment of these systems challenging.
@@ 8. Discussion
Through our systematic literature review and analysis of real-world CCS implementations, we have produced a framework that can help the HCI research community better reason about, and design, these systems. In this section, we discuss future directions for CCS research in HCI. Specifically, we focus on how CCS fit within the broader field of content moderation, highlight underexplored portions of the design space, and reflect on the normative implications of CCS design choices.
@@ 8.1. Can CCS Replace Content Moderation?
CCS were created with the stated goal of addressing the misinformation problem (Coleman, 2021), but it is unclear if the platforms implementing them are serious about this aim. For example, many see Meta’s choice to replace their U.S. fact-checking program with a CCS (Kaplan, 2025) as a thinly-veiled effort to curry political favor with a presidential administration critical of fact-checking (Nix et al., 2025). In fact, from a platform’s perspective, CCS have other appealing properties: they cost less than professional fact-checkers, they shield platforms from accusations of biased policy enforcement, and they offload the difficult responsibility of content moderation onto users. It may be expedient for platforms to build CCS in the current moment, but it is unclear what the role of such systems will be if the political climate shifts. Still, as long as CCS are in a position to exert influence over the information ecosystem, it is important for the HCI research community to study them. Most importantly, we emphasize the need to better understand how CCS perform compared to other content moderation systems. CCS may perform well in certain cases, but if on the whole they perform worse than the systems they are replacing, for example on the question of coverage, they should be used sparingly.
While it has been suggested that they can, CCS cannot effectively replace a fuller spectrum of content moderation tools, and were not designed to do so (Elliott, 2023). Content moderation is a broad term that encompasses a variety of different goals, including preventing crimes, mitigating lawful but awful behavior, and promoting high quality content (Grimmelmann, 2015). The tools used for content moderation are similarly varied and include both “hard” approaches, like content take-downs and account bans, and “soft” approaches, such as content labeling (Zannettou, 2021). Systems that provide context are narrow in both their goal and mechanism: they are a soft moderation attempt to aid content interpretation by providing users with additional information (Matamoros-Fernández and Jude, 2025). And yet, the empirical results summarized in this paper demonstrate that CCS may play a meaningful role in the content moderation ecosystem, and it is worth exploring how they may integrate with other parts of it. For example: how might CCS be integrated into platforms’ automated moderation systems? In the Platform Treatment design aspect, none of the implementations that we examined used the CCS output to affect the visibility of posts. Perhaps these systems can be integrated into platform content recommendations algorithms to promote or demote posts based on their notes.
@@ 8.2. Underexplored CCS Design Aspects
Our design space reveals several ways forward for a HCI research agenda on CCS. Several aspects are underexplored by current research, particularly the Participation aspect. Encouraging positive participation in collaborative computing systems is a classic topic in HCI research (Kraut and Resnick, 2012) and one in which the community could make strong contributions. One unexplored participation dimension is how platforms might more actively solicit notes and ratings from contributors. Past HCI work on how to best make use of the various skills of crowds may be instructive here (Bernstein et al., 2015; Kittur et al., 2013). It is worth noting that the more consequential these systems become, the more they will be targets for negative forms of participation as well, primarily gaming (Daniel et al., 2018; Epstein et al., 2020). Bridging algorithms are designed to be resistant to certain types of gaming (Wojcik et al., 2022), but vulnerabilities may emerge. One other area for study is the Presentation aspect, specifically how different approaches to visually displaying notes may affect user understanding of misleading claims (Gamage et al., 2025; Martel and Rand, 2023). Another promising area is the Inputs aspect, where recent work has begun to explore the ways in which LLMs can supplement human effort and increase the quality and quantity of notes (De et al., 2025; Mohammadi and Yasseri, 2025). Of particular interest is X’s recent decision to allow AI note contributors via a new API (Li et al., 2025). How such decisions affect the quantity and quality of notes should be a near-term focus of HCI research. Of course, studying how changes to CCS impact information dynamics across platforms will be difficult without changes in the Transparency aspect. Transparency about how CCS works (as X has provided to date) facilitates greater research insights. Hopefully other platforms will soon adopt similar approaches to support ongoing studies of CCS effectiveness.
@@ 8.3. Normative Guidelines for CCS Design
CCS are complex sociotechnical systems (Cherns, 1976) that integrate people, algorithms, and speech. These systems’ growing influence over the information shown on major social media platforms means that the design choices made in their implementation can have significant social consequences (Friedman and Nissenbaum, 1996; Flanagan et al., 2008). With such high stakes, it is crucial that a human-centered research agenda consider the implications of different design choices, in order to holistically evaluate these systems’ impacts. One way to do so is to identify the primary normative concerns of such a system. We emphasize three interconnected concerns that should guide CCS design: user informedness, distribution of power, and fairness.